Find the Laplace transform of f(t)=e^-t,where f(t+1)=f(t)
1
Expert's answer
2020-05-21T16:28:24-0400

If "f(t+T)=f(t)," then

"F(s)=L(f(t))={\displaystyle\int_{0}^T e^{-st}f(t)dt\over 1-e^{-sT}}"

"f(t+1)=f(t),\ f(t)=e^{-t}"




"F(s)=L(f(t))={\displaystyle\int_{0}^1 e^{-st}e^{-t}dt\over 1-e^{-s(1)}}="

"=-{1\over (1-e^{-1})(s+1)}\bigg[e^{-(s+1)t}\bigg]\begin{matrix}
 1\\
 0
\end{matrix}="

"={1-e^{-(s+1)}\over (1-e^{-1})(s+1)}"